# Imprints of Supermassive Black Hole Evolution on the Spectral and Spatial Anisotropy of Nano-Hertz Stochastic Gravitational-Wave Background

### v1

@@ 4 Adaptive technique: Cosmological scales to sub-pc scale
Cosmological simulations, which aim to model the evolution of the large-scale distribution of the Universe, are among the most memory-intensive simulations in astrophysics. Simulating a $\rm Gpc^{3}$ comoving volume at kpc resolution can take a long time, making kpc-resolution simulations feasible only for small volumes. High resolution simulation captures detailed halo and galaxy dynamics, while large-scale simulations offer insights into the large-scale structure of the Universe. The small-scale dynamics do not affect the inference of the large-scale simulation due to the low resolution of these simulations. Therefore, it becomes essential to bridge the large-scale cosmological simulation to the small-scale simulations by modeling the impact of the physics at the small scale on the larger-scale structure for studying the signatures of the cosmic evolution of SMBHBs from the n-Hz SGWB signal.
@@ 4.1 Cosmological scales to galaxy scales
Within the domain of cosmological simulations, bridging the vast range of scales from the largest cosmological structures to individual galaxies presents a significant challenge. Simulations such as MICECAT and ROMULUS (Tremmel et al., 2017; Butsky et al., 2019; Tremmel et al., 2020; Saeedzadeh et al., 2023b) are at the forefront of this effort. MICECAT covers one-eighth of the sky covering around 4 $\rm Gpc^{3}$ of comoving volume. It focuses on the evolution of large-scale structures, tracing the growth of cosmic web-like filaments and the formation of dark matter particles over billions of years. On the other hand, the ROMULUS simulations are high-resolution, cosmological simulations that explore the smaller, more intricate scales of the Universe. Among the ROMULUS simulations, ROMULUS25 is the leading simulation that extends over a uniform volume of $\rm 25~{}Mpc^{3}$ (Saeedzadeh et al., 2023a; Sharma et al., 2020). The simulation includes phenomena such as star formation, feedback mechanisms, and interactions with the surrounding environment. It employed a massively parallel tree+SPH code, CHaNGa (Menon et al., 2015; Wadsley et al., 2017), and has the Plummer equivalent gravitational force softening of 250 pc, a maximum SPH resolution of 70 pc. This high resolution enables the tracking of the evolution of galaxies and the SMBHs to the sub-kpc scale. One of the key findings from the ROMULUS25 simulation is the insights into the properties of the SMBH population and the intricate relationship between SMBH mass and the properties of their host galaxies and halos (Ricarte et al., 2021; Saeedzadeh et al., 2023a; Tremmel et al., 2023). These simulations reveal that SMBHBs crucial for PTA are more likely to reside in galaxies with high stellar and halo masses but low star formation rates (SFR) that reisdes at the center of galaxy group (Saeedzadeh et al., 2023a). Specifically, SMBHBs with chirp masses greater than $10^{8}M_{\odot}$ are more prevalent in galaxies with stellar masses exceeding $10^{10.5}M_{\odot}$ (Saeedzadeh et al., 2023a). The characteristics of SMBHBs from a small-volume simulation of ROMULUS25 can be complemented by cross-matching them with MICECAT based on their stellar mass and halo mass. Using the knowledge gained from the ROMULUS simulations, we can effectively populate galaxies in MICECAT with SMBHBs using parametric relation between $M_{\rm BH}$ and $M_{*}$.
@@
It is to be noted here that the MICECAT catalog is complete for DES-like surveys ($i<24$) out to z=1.4 (DES, 2005). Therefore, the completeness of the catalog is poor at high redshifts. As a result, the number of galaxies in the catalog at high redshifts is absent. This limits us from exploring high redshift (beyond $z>1.4$) SMBHB mergers from the simulated Universe. However, in standard scenarios, the mergers of SMBHs that can contribute to the n-Hz signal take a significant fraction of the age of the Universe, to come from kpc to sub-pc scale (Saeedzadeh et al., 2023a). This leads to a larger delay time from the time SMBHs form to the time SMBHBs can contribute to the n-Hz frequency band. As a result, the number of SMBHs that can contribute to the n-Hz signal at high redshift is lower, as also shown in Saeedzadeh et al. (2023a) from ROMULUS25. From the simulation, we have seen that contributions to the n-Hz signal mostly arise from redshift $z<1.4$. Also, the masses of SMBHs are likely to be lighter at higher redshifts than at lower redshifts, as they have more cosmic time to gain masses through accretion, except for the scenario where massive BH seeds are formed by the direct collapse of massive gas clouds Natarajan et al. (2023). Consequently, the masses that can contribute to the n-Hz signal will be lighter than the sources at lower redshifts, making their contribution to the n-Hz signal weak. Due to both these effects, we expect the limitation of MICECAT due to the redshift cutoff not to be severe for our analysis.
@@ 5 Gravitational Wave from SMBHBs
SMBHBs emit GWs in the frequency range $10^{-9}-10^{-7}$ Hz detectable by PTA. The orbit of the SMBHBs in this frequency range is highly stable. Fig. 3 illustrates the evolution of the frequency of GW emitted by SMBHBs starting with three different frequencies. In particular, SMBHBs with a chirp mass $10^{9}M_{\odot}$ emitting at a frequency as high as $3\times 10^{-8}$ Hz demonstrate a minimal frequency deviation over the observational timescale of PTAs. This makes SMBHBs a fairly monochromatic source of GW.
The SGWB is the superposition of all the GW sources that are unresolved as an individual source. It is defined as the GW energy density per unit logarithmic frequency divided by the critical energy density of the Universe. The SGWB density can be expressed as a function of frequency $f$ and sky position $\hat{n}$ as (Phinney, 2001; Sesana et al., 2008; Christensen, 2018)
@@
×$\times\left[\frac{1}{1+z}\frac{d^{n+2}N(\hat{n})}{d\Theta_{n}dVdt_{r}}\right]\left[\frac{1+z}{4\pi d_{L}^{2}c}\right]$
@@
where $\frac{d^{n+2}N(\hat{n})}{d\Theta_{n}dVdt_{r}}$ is the number of events per unit comoving volume per unit cosmic time ($t_{r}$) at a sky direction $\hat{n}$, $\frac{dE_{\rm{gw}}(f,\hat{n})}{df_{r}}$ is the energy emitted by the source per unit source frame frequency ($\rm f_{r}$) and $\Theta_{n}$ = $\{\theta_{i}\}_{i=1}^{n}$ denotes the GW source parameters. The quantity $\frac{d^{3}N}{d\Theta_{n}dVdt}$ can be expressed as
$\frac{d^{n+2}N(\hat{n})}{d\Theta_{n}dVdt}=\frac{d^{n+2}N(\hat{n})}{d\Theta_{n}dVd\ln f_{r}}\times\frac{d\ln f_{r}}{dt_{r}},$ (3)
where $\frac{d^{3}N(\hat{n})}{d\Theta_{n}dVd\ln f_{r}}$ represents the comoving number density of the sources emitting in logarithmic frequency interval. Substituting eq. 6 into eq. 2, gives
@@
×$\times\bigg{[}\frac{d^{n+2}N(\hat{n})}{d\Theta_{n}dVd\ln f_{r}}\bigg{]}\bigg{[}\frac{1}{4\pi d_{L}^{2}c}\bigg{]}$
@@
The term $\frac{d\ln f_{r}}{dt_{r}}\frac{dE_{\rm{gw}}(\Theta_{n},\hat{n})}{d\ln f_{r}}$ is the flux of energy emitted by a binary emitting at frequency $f_{r}$
$\frac{dE_{gw}}{d\ln f_{r}}=\frac{(G\pi)^{2/3}M_{c}^{5/3}\times f_{r}^{2/3}}{3},$ (5)
where $M_{c}$ is the chirp mass of the SMBHB. The relative fluctuation in $\Omega_{\rm gw}(f)$ can be defined as,
$\Delta\Omega_{gw}(f,\hat{n})\equiv{\frac{\Omega_{gw}(f,\hat{n})-\overline{\Omega}_{gw}(f)}{\overline{\Omega}_{gw}(f)}},$ (6)
where $\hat{n}$ is the line-of-sight direction. $\Delta\Omega_{gw}(f,\hat{n})$ can be written in terms of spherical harmonics $Y_{\ell m}(\hat{n})$ as
@@
$C_{\ell}(f)$ is the angular power spectrum. It captures the angular fluctuation of the $\Delta\Omega_{\rm gw}(\rm f,\hat{n})$ at angular scale of $\Delta\theta=\pi/\ell$. While the monopole ($\ell=0$) describes the isotropic $\Omega_{gw}$, the anisotropy depends on the higher order moment ($\ell>0$).
@@ 6.1 SMBHB Population
We aim to connect the SMBHB host galaxy properties to the population of SMBHBs. The distribution of the source population in eq. 4 can then be written with the quantities discussed in the Sec. 3 as
$\frac{d^{5}N}{d\log(M_{*})d\log(M_{\rm BH})dqdVda}=$ $\frac{dN}{d\log(M_{*})}\times P(M_{\rm BH},q,a|M_{*},z)$ (9)
$=$ $\overbrace{\frac{dN}{d\log(M_{*})}}^{MICECAT}$
@@
here, the term $\frac{dN}{d\log(M_{*})}$ represents the number of host galaxies with stellar mass, $M_{*}$, in the catalog. The terms $P(M_{\rm BH}|M_{,}z)$, $P(q|M_{,}z)$, and $P(a|M_{*},z)$ represent the probability distributions of the primary mass ($M_{\rm BH}$), mass ratio ($q$), and binary separation ($a$), respectively. The $\frac{dN}{d\log(M_{*})}$ can be directly obtained from the MICECAT simulation. While the distributions $P(M_{\rm BH}|M_{,}z)$ and $P(q|M_{,}z)$ can be obtained from the ROMULUS simulation (Saeedzadeh et al., 2023a) and by using parametric forms discussed below, the distribution $P(a|M_{*},z)$ can be modeled by including several dissipative processes that are active at sub-pc scales. Although the SMBH mass can also depend on several other galactic properties, for simplicity, we consider only the stellar mass dependence of SMBH mass.
We model $P(M_{\rm BH}|M_{,}z)$ and $P(q|M_{,}z)$ as
@@
where $\mathcal{N}$ is the normalised Gaussian with standard deviation $\sigma_{m}$ and mean $\mathrm{Log}_{10}(M_{\mu})$ as (Reines & Volonteri, 2015; Habouzit et al., 2021; Muhamed Kozhikkal et al., 2023; Saeedzadeh et al., 2023a)
@@
In eq. 11 $\eta$, $\rho$, and $\nu$ are free parameters, where $\nu$ controls the redshift evolution of the relation. This parametric form will make it feasible to obtain the theoretical signal of $\Omega_{\rm GW}(f)$ and $C_{l}(f)$ for a vast range of astrophysical models efficiently, which can be incorporated in a Bayesian framework to constrain from n-Hz data in future.
In Fig. 4, we show the primary mass ($M_{\rm BH}$) distribution of the simulated SMBHBs using the above equations for different values of $\eta$ and $\rho$. The scenario with higher values of $\eta$ and $\rho$ favors higher mass black holes, as shown in the figure. In Fig. 4, we display the primary mass distribution of the SMBHBs for different $\nu$ values at various redshift bins. A negative (or positive) value of $\nu$ indicates a preference for higher mass at low (or high) redshifts. At low redshifts, the distributions for all values of $\nu$ considered here appear to be very similar. However, at high redshifts, the distribution varies significantly for different $\nu$ values. This allows us to see the redshift evolution of SMBH masses in the Universe. From standard cosmological scenarios, we expect the heavier SMBHs to be more likely in the low redshifts, which is captured by $\nu<0$.
Figure 4: (a) Mass distribution of the primary mass $M_{\rm BH}$ of SMBHBs for different $\eta$ and $\rho$. (b) Mass distribution of the primary mass $M_{\rm BH}$ of SMBHBs for different $\nu$ at different redshift bins.
In Fig. 5, we show the primary mass of the SMBHs ($M_{\rm BH}$) and the stellar mass ($M_{*}$) of the host galaxy of the simulations, with their redshift represented in color-bar for different values of $\eta$ and $\rho$, and $\nu$. In Fig. 5, 5, and 5, $\nu$= 0, therefore SMBH mass-Stellar mass ($M_{\rm BH}-M_{*}$) relation does not evolve with the redshift. In Fig. 5 and 5, $\nu$= 0.5 and -0.5 respectively. The value $\nu=0.5$ indicates that there are heavier BHs at higher redshifts, while $\nu=-0.5$ indicates that there are lighter BHs at higher redshifts. This trend can be observed in the two figures, where most of the low-redshift BHs lie below and above the fitted lines in Fig. 5 and 5, respectively. If heavier SMBHs are more prevalent at lower redshifts, it would suggest that SMBHs are continuously growing even at low redshifts. However, if there is a population of SMBHs with heavier BHs at higher redshifts than at lower redshifts, this could imply that the growth of heavier SMBHs ceased at lower redshifts. This parameter $\nu$ also has a significant impact on the SGWB anisotropy signal as we will discuss in the results section.
@@
where $\frac{dN_{a}}{dt_{c}}$ and $\frac{dt_{c}}{da}$ represent the coalescence rate and the residence time of the binaries at the separation of ’$a$’, respectively. We have assumed a power law form of $\frac{dN_{a}}{dt_{c}}$, where $\delta$ is a power law index and $\mu$ controls its redshift evolution of the form
@@
The binary separation $a$ corresponds directly to the emitted frequency ($f_{r}$) of the binary, which is an observable quantity of gravitational waves in terms of redshifted frequency $f_{z}=f_{r}/(1+z)$ for a source contributing from redshift $z$. Thus, we prefer to model the separation distribution $P(a|M_{*})$ with the probability distribution of frequencies emitted by the binary $P(f_{r}|M_{*})$ as
@@
where $\frac{da}{df_{r}}$ is the Jacobian of the transformation from ’$a$’ to ’$f_{r}$’. We can write the number of coalescing binaries as a function of frequency $\frac{\mathrm{d}N_{f}}{\mathrm{d}t_{c}}$ as
@@
where the parameters $\delta_{f}$, $z_{0}$, and $\mu$ control the frequency dependence and the redshift dependence of the coalescing rate across cosmic time. The residence time for the GW signal to emit at a frequency $f_{r}$ $\frac{\mathrm{d}t_{c}}{\mathrm{d}f_{r}}$ is modeled as (Saeedzadeh et al., 2023a)
$\frac{\mathrm{d}t_{c}}{\mathrm{d}f_{r}}\propto f_{r}^{-11/3}\times\frac{\tau_{h}}{\tau_{GW}}(f_{r}).$ (17)
@@
As discussed in the Sec. 3, the quantity $\frac{\tau_{h}}{\tau_{GW}}(f_{r})$ represents the environment hardening time scale and the factor $f_{r}^{-11/3}$ captures the GW only hardening time scale. The residence time of the binary due to only GW emission at frequency $f_{r}$ is proportional to $f_{r}^{-11/3}$. The factor also represents the frequency distribution of the SMBHB population in the absence of the environmental effect and the frequency-dependent coalescence rate.
@@
In Fig. 6, we illustrate the $\frac{dN_{f}}{df}$ vs $f$, of the sources in the simulation along with the redshift represented by colormap as well as theoretical curves at three different redshifts, for different values of $\mu$ and $z_{0}$. In Fig. 6 and Fig. 6 we represent the case where $\mu$ = 0 and hence no redshift dependence of the frequency distribution, as can be seen from the analytical curves. In Fig. 6 and Fig. 6, we show the same quantity for the $\mu$ = 1, and $z_{0}$ = 0.5 and $z_{0}$ = 1, respectively. Here, the slope of the curve increases with the redshift. This means that there are comparatively more sources at low frequencies than at high frequencies at low redshifts. In these cases, the number of sources emitting at high frequency is even smaller than that of redshift independent case. Consequently, the SGWB power spectrum is expected to be less steep toward higher frequencies. Also, since the number of sources at high frequency is smaller, the anisotropy is expected to be larger in these cases.
@@
The analytical curves show the increasing trend in the slope with decreasing redshift. In this case, the slope of the curve is expected to be even larger than in the last case at low frequencies. The distribution of the sources at lower frequencies fits the analytical curve better than at the higher frequencies due to the greater number of GW sources at lower frequencies than higher frequencies. As the number of SMBHs must be an integer and cannot be fractional, there is a discrepancy between the analytical and simulated distributions of the sources at higher frequencies, where the analytical distribution predicts $(dN_{f}/df)\Delta f<1$. Due to the Poisson fluctuation on the SMBHB sources, the scatter of the sources at low frequencies is higher than at the lower frequencies.
@@
Figure 5: Plots showing the primary mass of the SMBHs ($M_{\rm BH}$) and the stellar mass ($M_{*}$) of the host galaxy of the simulations, with their redshift represented by colormap for different values of $\eta$ and $\rho$, and $\nu$.
@@
Figure 6: Plots showing the number sources emitting in each frequency bin ($\Delta f$) ($\frac{dN_{f}}{df}\times\Delta f$) vs Frequency ($f$), with their redshift (with bin size $\Delta z$ = 0.1) represented by colormap for different values of $\mu$, $z_{0}$, $\delta_{f}$, $\kappa$, and $\rho$.

### v2

@@ 3 Adaptive technique: Cosmological scales to sub-pc scale
Cosmological simulations, which aim to model the evolution of the large-scale distribution of the Universe, are among the most memory-intensive simulations in astrophysics. Simulating a $\rm Gpc^{3}$ comoving volume at kpc resolution can be very time-consuming, making kpc-resolution simulations feasible only for small volumes. High resolution simulation captures detailed halo and galaxy dynamics, while large-scale simulations offer insights into the large-scale structure of the Universe. The small-scale dynamics do not affect the inference of the large-scale simulation due to the low resolution of these simulations. Therefore, it becomes essential to bridge the large-scale cosmological simulation to the small-scale simulations by modeling the impact of the physics at the small scale on the larger-scale structure for studying the signatures of the cosmic evolution of SMBHBs from the n-Hz SGWB signal. A schematic diagram explaining the adaptive technique is shown in Fig. 1.
@@ 3.1 Cosmological scales to galaxy scales
Within the domain of cosmological simulations, bridging the vast range of scales from the largest cosmological structures to individual galaxies presents a significant challenge. Simulations such as MICECAT and ROMULUS (Tremmel et al., 2017; Butsky et al., 2019; Tremmel et al., 2020; Saeedzadeh et al., 2023b) are at the forefront of this effort. MICECAT covers one-eighth of the sky covering around 4 $\rm Gpc^{3}$ of comoving volume. It focuses on the evolution of large-scale structures, tracing the growth of cosmic web-like filaments and the formation of dark matter particles over billions of years. On the other hand, the ROMULUS simulations are high-resolution, cosmological simulations that explore the smaller, more intricate scales of the Universe. Among the ROMULUS simulations, ROMULUS25 is the leading simulation that extends over a uniform volume of $\rm 25~{}Mpc^{3}$ (Saeedzadeh et al., 2023a; Sharma et al., 2020). The simulation includes phenomena such as star formation, feedback mechanisms, and interactions with the surrounding environment. It employed a massively parallel tree+SPH code, CHaNGa (Menon et al., 2015; Wadsley et al., 2017), and has the Plummer equivalent gravitational force softening of 250 pc, a maximum SPH resolution of 70 pc. This high resolution enables the tracking of the evolution of galaxies and the SMBHs to the sub-kpc scale. One of the key findings from the ROMULUS25 simulation is the insights into the properties of the SMBH population and the intricate relationship between SMBH mass and the properties of their host galaxies and halos (Ricarte et al., 2021; Saeedzadeh et al., 2023a; Tremmel et al., 2023). These simulations reveal that SMBHBs crucial for PTA are more likely to reside in galaxies with high stellar and halo masses but low star formation rates (SFR) that reisdes at the center of galaxy group (Saeedzadeh et al., 2023a). Specifically, SMBHBs with chirp masses greater than $10^{8}M_{\odot}$ are more prevalent in galaxies with stellar masses exceeding $10^{11}M_{\odot}$ (Saeedzadeh et al., 2023a). The characteristics of SMBHBs from a small-volume simulation of ROMULUS25 can be complemented by cross-matching them with MICECAT galaxies. Using the knowledge gained from the ROMULUS simulations, we can effectively populate galaxies in MICECAT with SMBHBs using parametric relation between $M_{\rm BH}$ and $M_{*}$.
@@
It is to be noted here that the MICECAT catalog used in this work is complete for DES-like surveys ($i<24$) out to z=1.4 (DES, 2005). Therefore, the completeness of the catalog is poor at high redshifts. As a result, the number of galaxies in the catalog at high redshifts is absent. This limits us from exploring high redshift (beyond $z>1.4$) SMBHB mergers from the simulated Universe. However, in standard scenarios, the mergers of SMBHs that can contribute to the n-Hz signal take a significant fraction of the age of the Universe, to come from kpc to sub-pc scale (Saeedzadeh et al., 2023a). This leads to a larger delay time from the time SMBHs form to the time SMBHBs can contribute to the n-Hz frequency band. As a result, the number of SMBHs that can contribute to the n-Hz signal at high redshift is lower, as also shown in Saeedzadeh et al. (2023a) from ROMULUS25. From the simulation, we have seen that contributions to the n-Hz signal mostly arise from redshift $z<1.4$. Also, the masses of SMBHs are likely to be lighter at higher redshifts than at lower redshifts, as they have more cosmic time to gain masses through accretion, except for the scenario where massive BH seeds are formed by the direct collapse of massive gas clouds (Natarajan et al., 2023). Consequently, the masses that can contribute to the n-Hz signal at higher redshifts will be lighter than the sources at lower redshifts, making their contribution to the n-Hz signal weak. Due to both these effects, we expect the limitation of MICECAT due to the redshift cutoff not to be severe for our analysis.
@@ 4 Gravitational Wave from SMBHBs
SMBHBs can emit GWs in the frequency range $10^{-9}-10^{-7}$ Hz detectable by PTA. The orbit of the SMBHBs in this frequency range is highly stable. Fig. 2 illustrates the evolution of the frequency of GW emitted by SMBHBs starting with three different frequencies. In particular, SMBHBs with a chirp mass $10^{9}M_{\odot}$ emitting at a frequency as high as $3\times 10^{-8}$ Hz demonstrate a minimal frequency deviation over the observational timescale of PTAs. This makes SMBHBs a fairly monochromatic source of GW.
The SGWB is the superposition of all the GW sources that are unresolved as an individual source. It is defined as the GW energy density per unit logarithmic frequency divided by the critical energy density of the Universe ($\rho_{c}c^{2}$). The SGWB density due to the sources in a unit solid angle in the direction $\hat{n}$ can be expressed as a function of frequency $f$ as (Phinney, 2001; Sesana et al., 2008; Christensen, 2018)
@@
×$\times\left[\frac{1}{1+z}\frac{d^{n+4}N(z,\Theta_{n},\hat{n})}{d\Theta_{n}dVdt_{r}}\right]\left[\frac{1+z}{4\pi d_{L}^{2}c}\right]$
@@
where $V$ represents the comoving volume, $dV$ = $d\tilde{V}d\omega$, with $d\omega$ representing the solid angle element in the direction $\hat{n}$. The term $\frac{d^{n+4}N(z,\Theta_{n},\hat{n})}{d\Theta_{n}dVdt_{r}}$ is the mergers rate of the binaries per unit comoving volume at redshift z, at the direction $\hat{n}$, and per unit source properties, $\frac{dE_{\rm{gw}}(f,\Theta_{n},\hat{n})}{df_{r}}$ is the energy emitted by the source per unit source frame frequency ($\rm f_{r}$) and $\Theta_{n}$ = $\{\theta_{i}\}_{i=1}^{n}$ denotes the GW source parameters. The quantity $\frac{d^{n+4}N}{d\Theta_{n}dVdt_{r}}$ can be expressed as
$\frac{d^{n+4}N(z,\Theta_{n},\hat{n})}{d\Theta_{n}dVdt_{r}}=\frac{d^{n+4}N(z,\Theta_{n},\hat{n})}{d\Theta_{n}dVd\ln f_{r}}\times\frac{d\ln f_{r}}{dt_{r}},$ (3)
where $\frac{d^{n+4}N(z,\Theta_{n},\hat{n})}{d\Theta_{n}dVd\ln f_{r}}$ represents the comoving number density of the sources emitting in logarithmic frequency interval. Substituting eq. 3 into eq. 2, gives
@@
×$\times\bigg{[}\frac{d^{n+4}N(z,\Theta_{n},\hat{n})}{d\Theta_{n}dVd\ln f_{r}}\bigg{]}\bigg{[}\frac{1}{4\pi d_{L}^{2}c}\bigg{]}$
@@
The term $\frac{d\ln f_{r}}{dt_{r}}\frac{dE_{\rm{gw}}(\Theta_{n},\hat{n})}{d\ln f_{r}}$ is the flux of energy emitted by a binary emitting at frequency $f_{r}$ with
$\frac{dE_{\rm gw}}{d\ln f_{r}}=\frac{(G\pi)^{2/3}M_{c}^{5/3}\times f_{r}^{2/3}}{3},$ (5)
where $M_{c}$ is the chirp mass of the SMBHB. The relative fluctuation in $\Omega_{\rm gw}(f,\hat{n})$ can be defined as
$\Delta\Omega_{\rm gw}(f,\hat{n})\equiv{\frac{\Omega_{\rm gw}(f,\hat{n})-\overline{\Omega}_{\rm gw}(f)}{\overline{\Omega}_{\rm gw}(f)}},$ (6)
where $\Delta\Omega_{\rm gw}(f,\hat{n})$ can be written in terms of spherical harmonics $Y_{\ell m}(\hat{n})$ as
@@
$C_{\ell}(f)$ is the angular power spectrum. It captures the angular fluctuation of the $\Delta\Omega_{\rm gw}(\rm f,\hat{n})$ at angular scale of $\Delta\theta=\pi/\ell$. The anisotropy depends on the higher order moment ($\ell>0$). The relative fluctuations map $\Delta\Omega_{\rm gw}(f,\hat{n})$ is a mean zero field and is directly related to the tracer of the galaxy density field in the Universe (as discussed later in the paper). The mean-zero fluctuations map $\Delta\Omega_{\rm gw}(f,\hat{n})$ defined in this analysis is useful to recover an unbiased non-zero cross-correlation signal with the galaxy density field fluctuations map $\Delta\delta{\rm g}(\hat{n})$, in contrast to $\Omega_{\rm gw}(f,\hat{n})$ usually used in the literature (Agazie et al., 2023c; Pol et al., 2022).
@@ 5.1 SMBHB Population
We aim to connect the SMBHB host galaxy properties to the population of SMBHB. The distribution of the source population in eq. 4 can then be written with the quantities discussed in the Sec. 2 as
$\frac{d^{7}N}{d\log(M_{*})d\log(M_{\rm BH})dqdVda}\propto$ $\frac{d^{4}N}{d\log(M_{*})dV}P(M_{\rm BH},q,a|M_{*},z)$ (9)
$=$ $\overbrace{\frac{d^{4}N}{d\log(M_{*})dV}}^{MICECAT}$
@@
where, the term $\frac{d^{4}N}{d\log(M_{*})dV}$ represents the number density of host galaxies with stellar mass, $M_{*}$, in the catalog. The terms $P(M_{\rm BH}|M_{,}z)$, $P(q|M_{,}z)$, and $P(a|M_{*},z)$ represent the probability distributions of the primary mass ($M_{\rm BH}$), mass ratio ($q$), and binary separation ($a$), respectively, given the stellar mass of the host galaxy and its redshift. The $\frac{d^{4}N}{d\log(M_{*})dV}$ can be directly obtained from the MICECAT simulation. The distributions $P(M_{\rm BH}|M_{,}z)$ and $P(q|M_{,}z)$ can be estimated from the ROMULUS simulation (Saeedzadeh et al., 2023a) and by using parametric forms discussed below, the distribution $P(a|M_{*},z)$ can be modeled by including several dissipative processes that are active at sub-pc scales. Although the SMBH mass can also depend on several other galactic properties, for simplicity, we consider only the stellar mass dependence of the SMBHB masses. We model $P(M_{\rm BH}|M_{,}z)$ and $P(q|M_{,}z)$ as
@@
where $\mathcal{N}$ is the normalised Gaussian with standard deviation $\sigma_{m}$ and mean $\mathrm{Log}_{10}(M_{\mu})$, which is modeled as
@@
where $\eta$, $\rho$, and $\nu$ are free parameters, where $\nu$ controls the redshift evolution of the relation. This parametric form will make it feasible to obtain the theoretical signal of $\Omega_{\rm gw}(f)$ and $C_{\ell}(f)$ for a vast range of astrophysical models efficiently, which can be incorporated in a Bayesian framework to constrain from n-Hz data in future.
In Fig. 3, we show the primary mass ($M_{\rm BH}$) distribution of the simulated SMBHBs using the above equations for different values of $\eta$ and $\rho$. The scenario with higher values of $\eta$ and $\rho$ favors higher mass BHs, as shown in the figure. In Fig. 3, we display the primary mass distribution of the SMBHBs for different $\nu$ values at various redshift bins. A negative (or positive) value of $\nu$ indicates a preference for higher mass at low (or high) redshifts. At low redshifts, the distributions for all values of $\nu$ considered here appear to be very similar. However, at high redshifts, the distribution varies significantly for different $\nu$ values. This allows us to see the redshift evolution of SMBH masses in the Universe. From standard cosmological scenarios, we expect the heavier SMBHs to be more likely in the low redshifts, which is captured by $\nu<0$.
Figure 3: (a) Mass distribution of the primary mass ($M_{\rm BH}$) of SMBHBs for different $\eta$ and $\rho$. (b) Mass distribution of the primary mass ($M_{\rm BH}$) of SMBHBs for different $\nu$ at different redshift bins.
In Fig. 4, we show the primary mass of the SMBHs ($M_{\rm BH}$) and the stellar mass ($M_{*}$) of the host galaxy of the simulations, with their redshift represented in color-bar for different values of $\eta$ and $\rho$, and $\nu$. In Fig. 4, 4, and 4, $\nu$= 0, therefore the SMBH mass-Stellar mass ($M_{\rm BH}-M_{*}$) relation does not evolve with the redshift. In Fig. 4 and 4, $\nu$= 0.5 and $\nu$= -0.5 respectively. The value $\nu=0.5$ indicates that there are heavier BHs at higher redshifts, while $\nu=-0.5$ indicates that there are lighter BHs at higher redshifts. This trend can be observed in the two figures, where most of the low-redshift BHs lie below and above the fitted lines in Fig. 4 and 4, respectively. If heavier SMBHs are more prevalent at lower redshifts, it would suggest that SMBHs are continuously growing even at low redshifts. However, if there is a population of SMBHs with heavier BHs at higher redshifts than at lower redshifts, this could imply that the growth of heavier SMBHs ceased at lower redshifts. This parameter $\nu$ also has a significant impact on the SGWB anisotropy signal as we will discuss in the results section.
@@
where $\frac{dN_{a}}{dt_{c}}$ and $\frac{dt_{c}}{da}$ represent the coalescence rate and the residence time of the binaries at the separation of ’$a$’, respectively. We have assumed a power-law form of $\frac{dN_{a}}{dt_{c}}$, where $\delta$ is a power law index and $\mu$ controls its redshift evolution, given by:
@@
The binary separation ’$a$’ corresponds directly to the emitted frequency ($f_{r}$) of the binary, which is an observable quantity of gravitational waves in terms of redshifted frequency $f_{z}=f_{r}/(1+z)$ for a source contributing from redshift $z$. Thus, we prefer to model the separation distribution $P(a|M_{*})$ with the probability distribution of frequencies emitted by the binary $P(f_{r}|M_{*})$ as
@@
where $\frac{da}{df_{r}}$ is the Jacobian of the transformation from ’$a$’ to ’$f_{r}$’. We can write the coalescence rate as a function of frequency as:
@@
where the parameters $\delta_{f}$, $z_{0}$, and $\mu$ control the frequency dependence and the redshift dependence of the coalescing rate across cosmic time. The residence time of a binary at a frequency $f_{r}$, $\frac{\mathrm{d}t_{c}}{\mathrm{d}f_{r}}$, is modeled as (Saeedzadeh et al., 2023a)
$\frac{\mathrm{d}t_{c}}{\mathrm{d}f_{r}}\propto f_{r}^{-11/3}\times\frac{\tau_{h}}{\tau_{\rm GW}}(f_{r}).$ (17)
@@
As discussed in the Sec. 2, the quantity $\frac{\tau_{h}}{\tau_{\rm GW}}(f_{r})$ represents the environment hardening time scale and the factor $f_{r}^{-11/3}$ captures the GW only hardening time scale. The residence time of the binary due to only GW emission at frequency $f_{r}$ is proportional to $f_{r}^{-11/3}$. The factor also represents the frequency distribution of the SMBHB population in the absence of the environmental effect and the frequency-dependent coalescence rate.
@@
In Fig. 5, we illustrate the $\frac{dN_{f}}{df}$ vs $f$, of the sources in the simulation along with the redshift represented by colormap as well as theoretical curves at three different redshifts, for different values of $\mu$ and $z_{0}$. In Fig. 5 and Fig. 5 we represent the case where $\mu$ = 0 and hence no redshift dependence of the frequency distribution, as can be seen from the analytical curves. In Fig. 5 and Fig. 5, we show the same quantity for the $\mu$ = 1, with $z_{0}$ = 0.5 and $z_{0}$ = 1, respectively. Here, the slope of the curve decreases with the redshift (less negative). This means that there are comparatively more sources at low frequencies than at high frequencies at low redshifts. In these cases, the number of sources emitting at high frequency is even smaller than that of redshift independent case. Consequently, the SGWB power spectrum is expected to be less steep toward higher frequencies. Also, since the number of sources at high frequency is smaller, the anisotropy is expected to be larger in these cases.
@@
The analytical curves show the increasing trend in the slope with decreasing redshift. In this case, the slope of the curve is expected to be even larger than in the last case at low frequencies. The distribution of the sources at lower frequencies fits the analytical curve better than at the higher frequencies due to the larger number of sources at lower frequencies than higher frequencies. As the number of SMBHs must be an integer and cannot be fractional, there is a discrepancy between the analytical and simulated distributions of the sources at higher frequencies, where the analytical distribution predicts $(dN_{f}/df)\Delta f<1$. Due to the Poisson fluctuation in the number of sources, the scatter of the sources at high frequencies is higher than at the low frequencies.
@@
Figure 4: Plots showing the primary mass ($M_{\rm BH}$) of the SMBHBs and the stellar mass ($M_{*}$) of the host galaxies from the simulations, for different values of $\eta$ and $\rho$, and $\nu$, with their redshift represented by colormap
@@
Figure 5: Plots showing the number of sources emitting in each frequency bin $\Delta f$ = $2\times 10^{-9}$ Hz, and redshift bin $\Delta z$ = 0.1, as a function of the observer frame Frequency ($f$) for different $\mu$, $z_{0}$, $\delta_{f}$, and $\kappa$. The redshift is represented by a colormap.